The rate law for the decomposition of PH3 is, R = -d[PH3]/dt = k[PH3]. It takes 120 s for the concentration of 0.84 M PH3 to decrease to 0.21 M. How much time is required for 1.5 M PH3 to decrease to 0.56 M?
Answer #1

We are given the rate as R = -d[PH3]/dt = k[PH3] where [PH3] represents concentration of PH3. We can take [PH3] on the left-hand side and then find the integrated rate law.

The integrated rate law can be written as ln[PH3]t - ln[PH3]0 = -kt

Here [PH3]t and [PH3]0 represent the concentrations of PH3 at time t seconds and at time 0 seconds respectively.

From the integrated rate law, ln[PH3]t = ln[PH3]0 - kt

ln(0.21) = ln(0.84) - k(120)

k = 0.01155 sec-1

Substituting the value of rate constant k in the integrated rate law in order to find the required time,

ln(0.56) = ln(1.5) - (0.01155)t

t = 85.306 seconds